# Federated Foundation Model for Cardiac CT Imaging

### v1

@@ 2 Results
Figure 4: Comparison of locally and federated trained UNets and transformer-based models (SWIN-UNETR). The locally trained models perform well on their client’s respective data, but do not generalize to the data from other clients. The transformer-based architecture performs worse than the Unet. The generalization performance can be enhanced with federated training, but the UNet still performs and generalizes better. After performing federated KD and subsequent finetuning the performance of the transformer-based model is on par with the UNet on detecting the hinge points, coronary ostia, and membranous septum, while outperforming it on segmenting the calcification. While the predictive performance of the SWIN-UNETR can be enhanced with more training samples due to KD to be better or on par with the UNet architecture, KD does not enhance the performance of the UNet to a similar degree. Further, the SWIN-UNETR generalizes better to other downstream tasks than the UNet based architecture.
@@
One crucial aspect that hinders the widespread usage of federated learning to data is the inferior label quality sometimes present at participating locations. In the centralized setting one can identify false labels from training with visual inspection. Due to its inherent privacy constraints original data such as images cannot be shared and inspected, however, their annotations can be exchanged. We therefore compared the geometric relation of labels to each other across participating locations to find outliers or a systematic bias. For example, we identified a mixed up of label ids for upper and lower membranous septum point, shown in Fig. 5.
Fig. 5 shows some outliers for hinge points and membranous septum. Interestingly, the spread of labels is larger in the manual annotations, while the predictions of the network are more centered. This indicates a higher inter-observer variability, which we separately assessed in the following section. Furthermore, no confusion of point ids occurred in the predicted landmarks.
@@
To quantify the inter-observer variability of the manual generated ground truth to our model, we evaluated the performance of our final model on the public ImageCAS dataset [61] against each annotator from the participating locations, which has labeled 20 samples. The mean distance from the mean over all annotations is $2.60\pm 3.58\,\mathrm{mm}$. Using the same method for displaying the distribution of labels as in Fig. 5 the differences between human annotators from different hospitals are qualitatively explored. Despite providing a unified annotation protocol before labelling, some systematic biases can be found, e.g., between location 2 and 4 on the hinge point of the right coronary cusp (c.f. Fig. 6). For evaluation of the trained models the 2 mm pose a lower bound for the test error and our results show that our model is almost on-par (Fig. 6).
Figure 6: Distribution of hinge point labels across annotators and boxplot of distance of annotators and trained models on public dataset ImageCAS [61], which also serves as out-of-distribution testset. a) All annotators have placed the hinge points at the correct location. However, some systematic differences can be observed (e.g. between RCC of location 2 and 4). b) Average distance from mean points. The median distance for the human annotators is around 2 mm. Convolutional networks generalize better for local (on one location only) and federated training. By using KD on a large dataset, the performance and generalizability of transformers can be significatnly enhanced.
@@
Beyond planning of TAVI procedure, pre-procedural exclusion of relevant coronary artery disease is recommended in these patients by current guidelines [39]. To investigate the generalization performance of our trained models, we opt for segmenting the coronary arteries in the public ImageCAS dataset [61], which already includes contours of the vessel lumen for 1000 patients (80/20 train-test-split).
For segmenting the coronary arteries, we restrict ourselves to only finetune the last output layer of both models trained with KD, the UNet and SWIN-UNETR, to test the amount of meaningful features already extracted by the backbone of the federated model. In both models the last layer is a $1\times 1$ convolution that only reweighs the feature maps from the previous layer. While the SWIN-UNETR yields a DICE score of 24.45 the UNet is only able to achieve a DICE score of 4.45. We attribute this to the learning of global context in the transformer encoder that enables better performances in foundation models compared to convolutional based ones.
3 Discussion
We performed the largest federated learning cardiac imaging study to date on 8124 CT scans across eight hospitals in Germany. We are the first to solve the problem of federated learning on partially labeled clients in the realm of real world medical data instead of carefully curated public challenge datasets. In addition to training on labeled subsets of the data we also leverage the unlabeled images to increase the performance with semi-supervised federated knowledge distillation from a UNet teacher model to a transformer student model (SWIN-UNETR) [12]. The predictions of the federated trained submodels are better on the other locations compared to the single models trained on each location independently. Surprisingly, the federated model often performs better than the own local trained one. We attribute this to the better generalization ability of the federated model since our annotated training subsets are sometimes quite small and exhibit inter-observer variability. The federated workflow is especially beneficial for these locations that do not possess large quantities of (labeled) data. Our distilled SWIN-UNETR can serve as a foundation model for future work on cardiac CT imaging. Moreover, we have shown its generalizability for out of distribution samples on a publicly available dataset (see Fig. 6).
@@ 3 Discussion
Compared to other federated learning studies our work is of higher complexity [36, 9, 24] due to different field of views and anisotropic spacing. Contrary to past studies where all labels for all tasks existed at all clients we deal with partially labeled clients that have a skewed distribution of present labels. Approaches to learning from partially labeled datasets in a federated environment include learning one encoder per participating client and label [60]. However, this is only possible if each client is in possession of only one label. Further, marginal loss [46] is a popular method for dealing with partially labeled clients [26, 57]. The homogeneous distribution of anatomical structures in the human body can also be utilized in the training process to make assumptions about missing labels [62]. But the works are performed on large, relatively easy to segment structures (e.g. large organs such as liver). Different classification heads for each dataset in the training distribution also represent one way of dealing with partially labeled datasets [55]. However, this discards information from possibly intersecting labels across the datasets [54].
@@
Further, we are the first to employ knowledge distillation in a federated environment on real world data CT cardiac imaging data. Our final model that is distilled from three teacher models can perform the tasks of point detection and segmentation simultaneously. The problem solved in this work requires expert physician knowledge in contrast to solving a problem that only has a binary discriminative outcome that can be read out from a electronic health record database. The research on federated knowledge distillation (KD) shows similarities, as these studies are conducted using publicly available datasets [20]. In KD, the predictive ability of a low-capacity student network is enhanced by training it to align its predictions with those of a high-capacity teacher network [14]. Typically, knowledge is distilled from a group of teacher networks in FL, each trained on data from a different client [45, 31]. Other methods include distilling knowledge by matching attention maps between client models or aligning the feature maps of both models [10, 59, 52]. Wang et al. use marginal loss together with KD to learn a model across partially labeled clients. However, marginal loss was sufficient to learn all structures present in the federation in KD was employed to further enhance results [57]. As stated, this is only the case when trained on large labels that are relatively easy-to-segment.
@@
Before being able to train a model successful in a federation many tedious and practical obstacles needed to be solved. We were only allowed to initialize communication from within the clinic networks. Further, we had to take additional security measures in the form of transport layer security (TLS) and username and password authentication. We thus opted for fedbiomed as our library to enable federated learning, as they support many securtiy features out of the box [47]. Before even starting to gather the required data each location had to independently get an ethics agreement granted that allows training on that data. Although technically federated learning is an established paradigm for most legal departments the topic was new ground, which we have now covered with our study. We hope that the preprocessing and training scripts for this study can be used to accelerate further studies in the future.
@@
Once each location had successfully applied for the ethics agreement, the downloading of data from the PACS and other clinical information systems could be initiated. Although the system is standardized even the intra-hospital variance of data was large so that site specific pre-processing was necessary. Each hospital had different preferences regarding the recorded field of view and spacing. Different naming schemes made it difficult to extract the right series for each patient. Despite all the obstacles we believe one reason why our distilled model pretrained on the unlabeled data performs better is the large data heterogeneity induced by above factors.
@@
Federated learning has a privacy by design structure since no data leaves the individual hospitals. However, some works have proven that in a dishonest environment clients can either corrupt the training process or reconstruct part of model’s training data from the weights [56]. Multiple advanced privacy options exist that mitigate above points but come at the expense of model’s predictive performance. Examples include differential privacy [50, 1], secure multi-party computation, and homomorphic encryption [18]. However, we assume an honest environment where we must not be concerned about individual clients wanting to corrupt the training or inference procedure.
The model weights of the federated trained SWIN-UNETR foundation model are made available as a contribution to open science to enable further research in the cardiac CT heart imaging on more and diverse downstream tasks. The federated infrastructure is planned to be re-used for more use cases within the DZHK to enable large-scale AI in cardiovascular research. Concurrently, more hospitals are joining the federated network.
4 Methods

### v2

@@ Results
Figure 2: Comparison of UNets and transformer-based model (SWIN-UNETR) in boxplots for local, federated, and federated KD training for a) Hinge Points & Coronary Arteries (HPS & CAs), b) Memebranous Septum (MS), and c) Calcification. Test results on training clients are shown in blue, the results on independent test clients is shown in orange. In the boxplots median, 25th and 75th quartile, as well as outliers are shown. The locally trained models perform well on their locations’s respective data, but do not generalize to the data from other locations. The transformer-based architecture performs worse than the Unet. The generalization performance can be enhanced with federated training, but the UNet still performs and generalizes better. After performing federated KD and subsequent finetuning the performance of the transformer-based model is on par with the UNet on detecting the hinge points, coronary ostia, and membranous septum, while outperforming it on segmenting the calcification. While the predictive performance of the SWIN-UNETR can be enhanced with more training samples due to KD to be better or on par with the UNet architecture, KD does not enhance the performance of the UNet to a similar degree.
As can be seen in Figure 2a models trained only on the local data shards underperform on datasets from other locations. Transformer based architectures generalize worse than convolutional UNet based ones, which we attribute to the inherent inductive bias of these architectures. The mean distance of the predicted hinge points of the local UNet approach is at $3.09\pm 1.71\,\mathrm{mm}$ for the same location and at $3.80\pm 2.02\,\mathrm{mm}$ for held out test locations, while the SWIN-UNETR predicts points at a mean distance of $2.66\pm 1.79\,\mathrm{mm}$ and $4.89\pm 4.08\,\mathrm{mm}$ respectively. The ViT-based model overfits the training data significantly so that it even predicts points far off for the test sets on training clients ($18.43\pm 20.51\,\mathrm{mm}$ and $17.71\pm 19.42\,\mathrm{mm}$). Federated training improves generalization performance for both methods. However, the UNet ($2.59\pm 1.76\,\mathrm{mm}$, $3.43\pm 1.79\,\mathrm{mm}$) performs better than the SWIN-UNETR ($3.06\pm 1.70\,\mathrm{mm}$, $3.89\pm 1.91\,\mathrm{mm}$). While ViT can be improved with federated training its performance still falls short of the other two models ($5.97\pm 7.73\,\mathrm{mm}$, $6.32\pm 6.27\,\mathrm{mm}$). While the performance of the SWIN-UNETR can be enhanced by performing semi-supervised federated knowledge distillation from the federated trained UNet on the previously unlabeled data samples at all locations the performance of the KD UNet is similar to the federated one. The predicted points lie at a mean distance of $2.80\pm 1.71\,\mathrm{mm}$ for the training locations and $3.36\pm 1.83\,\mathrm{mm}$ for the held out test locations for the transformer and at $3.18\pm 1.92\,\mathrm{mm}$ and $3.83\pm 2.12\,\mathrm{mm}$ for the UNet. The performance for ViT can also be improved with our two-stage federated learning strategy, but the results again fall short ($5.76\pm 3.25\,\mathrm{mm}$, $5.81\pm 4.32\,\mathrm{mm}$).
Figure 3: Qualitative results of the predicted labels of FedKD SWIN-UNETR. The predictions of our final distilled model were inspected by two experienced cardiologists verifying that the points are placed within the anatomical variance present. RCC: right coronary cusp, LCC: left coronary cusp, NCC: non-coronary cusp, RCO: right coronary ostium, LCO: left coronary ostium, MS1: upper, and MS2: lower point of membranous septum, Myo: myocardium, LA: left atrium, LV: left ventricle, RA: right atrium, RV: right ventricle, PA: pulmonary artery.
Table 1: Results of model architectures for the different learning schemes and label types. Three model types are investigated: convolutional UNet, a vision transformer (ViT) for segmentation, and SWIN-UNETR, which uses a siding attention approach different to a conventional ViT. All architectures are trained locally per location (Local), federated across labeled subsets (Fed), and with our federated knowledge distillation (FedKD) approach. Results are reported for locations the model was trained at (Training) and tested at the remaining (Other). All values are presented as $\mathrm{mean}\pm\mathrm{std}$.
Training Scheme Model HPs & COs $\downarrow$ [mm] MS $\downarrow$ [mm] Calc $\uparrow$ [DICE]
Training Other Training Other Training Other
Local UNet $3.48\pm 2.77$ $4.27\pm 2.94$ $3.01\pm 1.84$ $4.30\pm 1.82$ $0.708\pm 0.103$ $0.644\pm 0.290$
ViT $9.45\pm 11.87$ $14.85\pm 16.35$ $6.86\pm 11.14$ $37.88\pm 33.10$ $0.644\pm 0.184$ $0.474\pm 0.275$
SWIN $2.66\pm 1.79$ $4.89\pm 4.08$ $3.96\pm 2.19$ $4.06\pm 2.16$ $0.709\pm 0.190$ $0.466\pm 0.265$
Fed UNet $2.91\pm 2.54$ $3.75\pm 2.38$ $3.27\pm 2.02$ $3.75\pm 1.96$ $0.495\pm 0.209$ $0.391\pm 0.212$
ViT $4.75\pm 4,17$ $3.71\pm 1.88$ $3.82\pm 2.50$ $5.32\pm 4.98$ $0.671\pm 0.191$ $0.636\pm 0.285$
SWIN $3.53\pm 2.82$ $3.98\pm 2.05$ $3.03\pm 1.85$ $3.30\pm 1.60$ $0.683\pm 0.202$ $0.692\pm 0.232$
FedKD UNet $3.54\pm 2.85$ $4.25\pm 2.94$ $2.99\pm 1.81$ $3.40\pm 1.56$ $0.527\pm 0.209$ $0.526\pm 0.228$
ViT $4.70\pm 4.14$ $3.72\pm 1.88$ $3.28\pm 2.31$ $4.35\pm 2.34$ $0.562\pm 0.200$ $0.566\pm 0.240$
SWIN $3.04\pm 2.34$ $3.54\pm 2.12$ $2.95\pm 1.72$ $3.29\pm 1.45$ $0.646\pm 0.208$ $0.670\pm 0.231$
HPs & COs = Hinge Points and Coronary Ostia, MS = Membranous Septum, Calc = Calcification, KD = Knowledge Distillation
The performance for detecting the membranous septum is similar to localizing the hinge points. The UNet generalizes better with fewer data samples, but the SWIN-UNETR can be improved with semi-supervised federated knowledge distillation to even surpass the UNet on the unseen test locations. The local UNet predicts a mean distance of $3.01\pm 1.84\,\mathrm{mm}$ on the same client and $4.30\pm 1.82\,\mathrm{mm}$ on others, the local SWIN-UNETR predicts points at $3.96\pm 2.19\,\mathrm{mm}$ and $4.06\pm 2.16\,\mathrm{mm}$ distance respectively. The conventional ViT again overfits drastically to the training data distributions predicting points at $6.86\pm 11.14\,\mathrm{mm}$ for the testsets of the training clients and $37.88\pm 33.10\,\mathrm{mm}$ When training both models in a federated manner the SWIN-UNETR generalizes better when knowledge distillation is employed, what is seen with the lower standard deviation. The UNet’s mean distance lies at $3.40\pm 1.56\,\mathrm{mm}$, while the SWIN-UNETR’s is at $3.29\pm 1.45\,\mathrm{mm}$. The performance on the training locations is very similar (UNet: $2.99\pm 1.81\,\mathrm{mm}$, KDT: $2.95\pm 1.72\,\mathrm{mm}$). While the ViT consistently underperforms both models we can see a strong improvement by employing our two-stage learning strategy.
Segmenting the calcification in the aortic root leads to different results than the previous tasks. Both transformers perform better than the UNet especially when trained in a federated manner. The SWIN-UNETR reaches a DICE score of $0.683\pm 0.201$ on the testsets of the training locations and $0.692\pm 0.232$ on the held out test location, the ViT $0.671\pm 0.191$ and $0.636\pm 0.285$ respectively. The federated UNet only achieves a DICE score of $0.410\pm 0.209$ and $0.391\pm 0.212$. The model trained with KD is almost on par with the federated trained SWIN-UNETR with DICE scores of $0.646\pm 0.208$ and $0.670\pm 0.231$. We attribute the slightly worse performance to the concurrent point detection which seems to favor partly other image features than calcification segmentation. The ViT’s performance again falls short of the SWIN-UNETR, it achieves DICE scores of $0.562\pm 0.200$ and $0.566\pm 0.240$ respectively.
In conclusion, semi-supervised federated knowledge distillation enhances the predictive performance of a transformer based architecture (SWIN-UNETR) [43] to be better or to keep astride with the UNet based counter part. Further, the tasks of locating the hinge points, coronary ostia, and membranous septum as well as segmenting the calcification of the aortic root can be solved with one model despite the distributed label classes across different classes. While the two-stage learning strategy improves performance also for other transformer architectures such as the conventional ViT, the performance when employing a shifted window self-attention is better. As we have shown that SWIN-UNETR outperforms the ViT consistently, we will focus on that architecture in the following.
@@
One crucial aspect that hinders the widespread usage of federated learning to data is the inferior label quality sometimes present at participating locations. In the centralized setting one can identify false labels from training with visual inspection. Due to its inherent privacy constraints original data such as images cannot be shared and inspected, however, their annotations can be exchanged. We therefore compared the geometric relation of labels to each other across participating locations to find outliers or a systematic bias. For example, we identified a mix up of label ids for upper and lower membranous septum point, shown in Figure 4c.
Figure 4 shows some outliers for hinge points and membranous septum. Interestingly, the spread of labels is larger in the manual annotations, while the predictions of the network are more centered. This indicates a higher inter-observer variability, which we separately assessed in the following section. Furthermore, no confusion of point ids occurred in the predicted landmarks.
@@
To quantify the inter-observer variability of the manual generated ground truth to our model, we evaluated the performance of our final model on the public ImageCAS dataset [31] against each annotator from the participating locations, each of whom labeled 20 samples. The mean distance from the mean over all annotations is $2.60\pm 3.58\,\mathrm{mm}$. Using the same method for displaying the distribution of labels as in Figure 4a and 4c the differences between human annotators from different hospitals are qualitatively explored. Despite providing a unified annotation protocol before labelling, some systematic biases can be found, e.g., between location 2 and 4 on the hinge point of the right coronary cusp (c.f. Figure 5a). For evaluation of the trained models the 2 mm pose a lower bound for the test error and our results show that our model is almost on-par (Figure 5b).
@@
Figure 5: Distribution of hinge point labels across annotators and boxplot of distance of annotators and trained models on public dataset. Labels were obtained for the public dataset ImageCAS [31], which also serves as out-of-distribution testset. a) All annotators have placed the hinge points at the correct location. However, some systematic differences can be observed (e.g. between RCC of location 2 and 4). b) Average distance from mean points in boxplot (median as well as 25th and 75th quartile). The median distance for the human annotators is around 2 mm. Convolutional networks generalize better for local (on one location only) and federated training. By using KD on a large dataset, the performance and generalizability of transformers can be significantly enhanced.
@@
Besides planning of TAVI procedures, pre-procedural exclusion of relevant coronary artery disease is recommended in these patients by current guidelines [44]. To investigate the generalization performance of our trained models, we opt for segmenting the coronary arteries in the public ImageCAS dataset [31], which already includes contours of the vessel lumen for 1000 patients (80/20 train-test-split).
For segmenting the coronary arteries, we restrict ourselves to only finetune the last output layer of both models trained with KD, the UNet and SWIN-UNETR, to assess the extent of feature extraction already achieved by the backbone of the federated model. In both models the last layer is a $1\times 1$ convolution that only reweighs the feature maps from the previous layer. While the SWIN-UNETR yields a DICE score of 0.245 the UNet is only able to achieve a DICE score of 0.045. We attribute this to the learning of global context in the transformer encoder that enables better performances compared to convolutional based ones.
Discussion
We performed the largest federated learning 3D cardiac CT imaging study to date on 8104 scans across eight hospitals in Germany. We are the first to solve the problem of federated learning on partially labeled datasets in the realm of real world medical data instead of carefully curated public challenge datasets. In addition to training on labeled subsets of the data we also leverage the unlabeled images to increase the performance with semi-supervised federated knowledge distillation from a UNet teacher model to a transformer student model (SWIN-UNETR and ViT based) [42, 43]. The predictions of the federated trained submodels are better on the other locations compared to the single models trained on each location independently. Surprisingly, the federated model often performs better than the own local trained one. We attribute this to the better generalization ability of the federated model since our annotated training subsets are sometimes quite small and exhibit inter-observer variability. The federated workflow is especially beneficial for these locations that do not possess large quantities of (labeled) data. Our distilled SWIN-UNETR can serve as a base model for future work on cardiac CT imaging. Moreover, we have shown its generalizability for out of distribution samples on a publicly available dataset (see Figure 5). While both transformer-based architectures achive better performance with larger dataset sizes, the performance is better for the SWIN-UNETR, which uses shifted window self-attention. While ViT’s performance is also enhanced, it’s performance falls short. We leave it for future work to examine whether this difference might potentially be mitigated with even larger dataset sizes.
@@ Discussion
Compared to other federated learning studies our work is of higher complexity [9, 12, 45] due to different field of views and anisotropic spacing. Contrary to past studies where all labels for all tasks existed at all locations we deal with partially labeled ones that have a skewed distribution of present labels. Approaches to learning from partially labeled datasets in a federated environment include learning one encoder per participating client and label [46]. However, this is only possible if each client is in possession of only one label. Further, marginal loss [47] is a popular method for dealing with partially labeled datasets [48, 49]. The homogeneous distribution of anatomical structures in the human body can also be utilized in the training process to make assumptions about missing labels [50]. But the works are performed on large, relatively easy to segment structures (e.g. large organs such as liver). Different classification heads for each dataset in the training distribution also represent one way of dealing with partially labeled datasets [20]. However, this discards information from possibly intersecting labels across the datasets [51].
@@
Further, we are the first to employ knowledge distillation in a federated environment on real world data CT cardiac imaging data. Our final model that is distilled from three teacher models can perform the tasks of point detection and segmentation simultaneously. The problem solved in this work requires expert physician knowledge in contrast to solving a problem that only has a binary discriminative outcome that can be read out from a electronic health record database. The research on federated knowledge distillation (KD) shows similarities, as these studies are conducted using publicly available datasets [19]. In KD, the predictive ability of a low-capacity student network is enhanced by training it to align its predictions with those of a high-capacity teacher network [24]. Typically, knowledge is distilled from a group of teacher networks in FL, each trained on data from a different location [52, 53]. Other methods include distilling knowledge by matching attention maps between client models or aligning the feature maps of both models [23, 54, 55]. Wang et al. use marginal loss together with KD to learn a model across partially labeled datasets [49]. However, marginal loss was sufficient to learn all structures present in the federation and KD was employed to further enhance results. As stated, this is only the case when trained on large labels that are relatively easy-to-segment.
@@
Before being able to train a model successful in a federation many tedious and practical obstacles needed to be solved. We were only allowed to initialize communication from within the clinic networks. Further, we had to take additional security measures in the form of transport layer security (TLS) and username and password authentication. We thus chose fedbiomed as library for federated learning, as they support many securtiy features out of the box [56]. We hope that the preprocessing and training scripts for this study can be used to accelerate further studies in the future.
@@
Once each location had successfully applied for the ethics agreement, the downloading of data from the PACS and other clinical information systems could be initiated. Although the system is standardized even the intra-hospital variance of data was large so that site specific pre-processing was necessary. Each hospital had different preferences regarding the recorded field of view and spacing. Different naming schemes made it difficult to extract the right series for each patient. Despite all the obstacles we believe one reason why our distilled model pretrained on the unlabeled data performs better is the large data heterogeneity induced by some of above factors.
@@
Federated learning has a privacy by design structure since no data leaves the individual hospitals. However, some works have proven that in a dishonest environment clients can either corrupt the training process or reconstruct part of model’s training data from the weights [59]. Multiple attack vectors exist that can mostly be divided in privacy- and utility-centered attacks. Privacy-centered attacks describe the obtaining of information through unintentional leakages during training. One such attack is Model Inversion, in which an attacker might obtain data, which was used for training, by studying how specific inputs affect the model’s output [60, 61]. Other attacks include Attribute inference [62, 63], obtaining attributes of clients rather than data, and Membership Inference Attacks [64, 65], which allow the attacker to infer whether an individual was part of the training dataset.
In our experimental setup we assume an honest environment. Our consortium comprises locations that are familiar with one another and share a common goal of advancing research, while adhering to strict privacy constraints. In such scenarios FL offers a privacy by design structure during the learning process. Still, one must be concerned about attacks that can be carried out on the final resulting model weights such as Model Inversion. In previous work, multiple factors mitigating the possible success of such attacks have been published [59]. These include knowledge distillation, as the model learns a distribution over the proxy model’s output [66]. Less overfitting causes more general gradients that are less bound to individual samples in the training data set and, thus, complicate reconstruction of input data [59]. Employing regularization as well as larger batch sizes during training positively influence the privacy guarantees of models [6].
Cryptographic methods, such as differential privacy (DP), secure multi-party communication (SMPC), and homomorphic encryption (HE), can further enhance privacy guarantees of federated learning [5]. Differential privacy perturbs the gradient update or the input data with zero mean noise equipping each data samples with a plausible deniability of membership in the dataset [67, 68]. SMPC protects the model training and update process by distributing computations among multiple parties using secret sharing, while HE encrypts the input data, enabling computations to be performed directly on the encrypted inputs [56, 69]. However, applying cryptographic methods is out of the scope for this paper and we leave it to future work to investigate its influence. Further, we assured privacy guarantees with knowledge distillation and less overfitting due to large dataset sizes and regularization.
The model weights of the federated knowledge distilled SWIN-UNETR model are made available as a contribution to open science to enable further research in the cardiac CT imaging on more and diverse downstream tasks. The federated infrastructure is planned to be re-used for more use cases within the DZHK to enable large-scale AI in cardiovascular research. Concurrently, more hospitals are joining the federated network.
Methods